Week of 160606

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:
  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes

ALICE ATLAS CMS LHCB
  BNL FNAL  

Monday

Attendance:

  • local: Luca (SCOD+Storage), Maria A. (MW), Maria D. (GGUS), Eva (DB), Andrea (FTS), Ivan (ATLAS), Marcelo (LHCb), Marian (Network+Monitoring)
  • remote: Eric (BNL), Francesco (CNAF), Rolf (IN2P3), Andrew (NLT1), David (FNAL), Vincenzo (EGI), Tiju (RAL), Di Qing (TRIUMF), Ulf (offline, NDGF)

Experiments round table:

  • ATLAS reports (raw view) -
    • Activities:
      • MC16 simulation is being tested in production system. A new round of derivation production started today towards ICHEP, network traffic is increased due to this.
      • Review of derivations volume continues, above budget by 50%, semi-automated way of running on the common MC15c background samples (e.g. different pile-up profiles covered automatically) contributed to large volume, deletion request has been made to physics groups for unused samples.
      • Preparing for the ADC TIM meeting at CERN next week, June 15-17.
    • Problems:
      • Long standing issue with Condor/Cream causing CERN resources not fully used (https://its.cern.ch/jira/browse/ASPDA-196), patched binary provided by Condor team was deployed on the relevant CEs last week.
      • CERN-PROD_SCRATCHDISK was full, one user requested 100TB, deletion rate was slow due to small files (sorting by age), the queue was rearranged sorting by file size to speed up the deletion rate.

  • CMS reports (raw view) -
    • Sorry for the delayed report - it's actually not that much
    • CMS taking data high good rate
    • KIT had an outage of their SRM node due to hardware issues last week
    • Operational issue with CMS DAS (Data Aggregation Service)
      • Kept experts and web operators busy over the weekend
      • Some central services affected
      • Situation has improved now, real source still under investigation

  • ALICE -

  • LHCb reports (raw view) -
    • Activity
      • Currently mostly simulation jobs and reconstruction on the grid
    • Site Issues
      • T0: Data transfer from pit to offline (castor) did not work from 26 May - 31 May due to problems on LHCb side. Transfers fine now.
      • T1: SARA : tape system downtime from 30/05/2016 to 02/06/2016.
      • T1: RAL : tape systems fixed

Sites / Services round table:

  • ASGC:
  • BNL: NTR
  • CNAF:gpfs update completed accordingly to the EGI suggestion (security). Intervention on CMS storage to be discussed with the experiment (48 hours down). Intervention on the Bari nodes to fix a problem on the disk cache (this week), no impact on the experiments activity.
  • FNAL: NTR
  • GridPP:
  • IN2P3: Reminder of the already announced downtime next week, June 14th: combined maintenance of power supply, network, various hard/software; dCache and batch down all day (batch draining already during the preceding night), other services for shorter periods of time. See downtime declarations for details. NOTE that the EGI operations portal will be down in the morning,too; this impacts the DOWNTIME NOTIFICATIONS.
  • JINR: HC dropped to 80% at 01-06 due to batch system stuck for hours, running jobs are not affected. Investigating.
  • KISTI:
  • KIT:
  • NDGF:
    • Last week upgrade to dCache 2.16.0. No big surprises, everything is working.
    • There will be a 1h downtime at Monday 0700 UTC for one hour to upgrade the database (GOCDB will be created later today). Actual upgrade takes around 5 minutes, but we want a full backup first and it takes time.
  • NL-T1:
    • NTR for Nikhef
    • Networking issue from SARA to other Tier-1s (ongoing, started investigation after problem with the link to BNL)
  • NRC-KI:
  • OSG:
  • PIC:
  • RAL:
    • There was an outage of the tape system during the second part of last week. Some components inside the tape library were replaced on Wednesday (1st June ). However, there was a short circuit introduced by a faulty component and the library was taken down from Thursday (2nd) to Friday (3rd) for electrical checks to be carried out. These were completed Friday afternoon and the library hardware was running successfully from the end of the afternoon. However, there have been problems with the library control software crashing. We have an intervention tomorrow (7th June) announced in the GOC DB for us to investigate this.
  • TRIUMF: A 4-hour downtime is scheduled to upgrade dCache to version 2.13.31 on Wednesday, the downtime will start 17:00 (UTC time)

  • CERN computing services:
  • CERN storage services:
    • FTS: During the weekend there has been an issue with Overloaded VMs, which caused ~10% transfer failing (OTG:0030871). The problem has affected all VOs and LHCb reported it via a GGUS:121954.
    • FTS upgrade to v 3.4.6 today from 13:30 to 14:30 CEST. The new version fixes a small issue with excessive logging discovered last week after the upgrade to 3.4.5
    • CASTOR update 7th June from 14:00 to 16:00 (Public) and 8th June 10:00 to 13:00 (ALICE, CMS, LHCb)
    • EOS update for ATLAS in standby in coordination with ATLAS Tier-0 team
  • CERN databases: Reminder - OS updates this week (during TS). Schedule available in the IT SSB.
  • GGUS: The release went well. The FNAL test ALARM was closed today as it turned out that there was an issue with the configuration of their list server and it took some time to fix this. All tests here.
  • MW Officer:
  • Security:
  • Monitoring:
    • ETF will be updated next Tue 14th June
  • Networks:
    • GGUS:121687 RAL consistent loss - waiting for an upgrade of the router at RAL
    • GGUS:121905 BNL to SARA - throughput dropped between CERN and SARA sometime after the intervention on the ESNet router at CERN (GGUS:120957) - waiting for SARA to fix perfSONAR configuration

AOB:

Edit | Attach | Watch | Print version | History: r21 < r20 < r19 < r18 < r17 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r21 - 2016-06-07 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback